# The emergent dynamics of double-folded randomly branching ring polymers

### v1

@@ VI Summary and Conclusion
Specifically, we have reimplemented and generalized our earlier elastic lattice model for tightly double-folded ring polymers [24]. To this end, we have extended the model such that sets of compatible ring and tree degrees of freedom are now simultaneously updated and defined at all moments of time. Then, we have introduced tree node moves complementing the original ring monomer moves. In addition to the spontaneous creation and deletion of side branches and the diffusive local mass transport along the tree, the present extension allows for a spatial diffusion of the tree that preserves the secondary structure of the double-folded rings.
@@
We have explored the dynamics of tree-like polymer structures for three paradigmatic types of double-folded ring systems: ideal rings, isolated self-avoiding rings, and rings in the melt state. For the corresponding (combinations of) local moves, we have observed the expected signatures of the slithering of rings around the tree arising from the local transport of reptons, the ring-in-an-array-of-obstacles dynamics emerging from the additional possibility to absorb and grow side branches, and a Rouse-like dynamics due to the diffusion of tree nodes. In this respect, the dynamics of quenched trees in the melt state is particularly interesting. While the monomer diffusion is in almost perfect agreement with the expectation $g_{1}\sim t^{0.4}$ from the Rouse model for crumpled rings, there are also clear indications for a slow-down of the center-of-mass dynamics on intermediate scales. In particular, we have observed and analyzed a non-trivial dynamics emerging from the coupling of our ring monomer and tree node moves, which appears to asymptotically dominate the ring dynamics. At least qualitatively, the observed scaling of $\Delta D_{AB}\sim N^{-1}$ is compatible with the interpretation that the constraint release due to the annealing of the tree structure can induce a Rouse-like dynamics.
We note that numerous experimental investigations using fluorescent microscopy have explored the 3D dynamics of chromatin in cell nuclei of mammals [49, 50, 51, 52], drosophila [53, 54] , and yeast [55]. Additionally, studies on genome dynamics within bacterial nucleoid have been conducted [56, 57, 58, 59]. Many experiments reveal a subdiffusive behavior characterized by an exponent close to $0.4$: $g_{1}\sim t^{0.4}$, which are compatible with the behavior of the present model. However, observations [60, 61] of exponents in the range of $0.5-0.6$ beyond the entanglement or branching scale call for different explanations.
VII Simulation Software
@@
The simulation software is an open-source package, distributed under the GNU General Public License v3.03.03.03.0 and written in C++. It is available at [62]. All simulations were performed on CBPSmn computer cluster of ENS-Lyon.
VIII Acknowledgement
@@
We gratefully acknowledge the support of the Centre Blaise Pascal’s IT test platform at ENS de Lyon (Lyon, France) for facilities. The platform operates the SIDUS solution [63] developed by Emmanuel Quemener. We thank Angelo Rosa and Cedric Vaillant for valuable discussions, and Ali Farnudi for his technical support. IJ and RE acknowledge funding from CNRS 80 Prime (MIMIC project).
@@
Appendix A Anpproximate solution of the null model for the dynamics resulting from independent additive contributions
Consider two dynamics, $A$ and $B$, generating MSDs $g_{A}(t)$ and $g_{B}(t)$, respectively. We are interested in the MSD $g_{AB}(t)$ of mixed schemes, where the local degrees of freedom are updated by $A$ and $B$ moves with probabilities $p_{A}$ and $p_{B}$, respectively, and where typically (but not necessarily) $p_{A}+p_{B}=1$. Our objective below is to estimate the full time dependence of the emergent behavior for the null model of independent and additive contributions from the $A$ and $B$ dynamics.
On first sight, the proper generalization of Eq. (11) would seem to be
$g_{AB}(t)=p_{A}g_{A}(t)+p_{B}g_{B}(t).$ (17)
However, with $\lim_{p_{B}\rightarrow 0}g_{AB}(t)=p_{A}g_{A}(t)$ this form in general does not reproduce the exactly known limit of $\lim_{p_{B}\rightarrow 0}g_{AB}(t)=g_{A}(p_{A}t)$: a pure scheme, where only a fraction $p_{A}$ of moves is accepted, has no other effect than to reduce the effectively elapsed time by a corresponding factor. The ansatz
$g_{AB}(t)=g_{A}(p_{A}t)+g_{B}(p_{B}t)$ (18)
overcomes this limitation. But while it reduces to Eq. (11) for simple diffusion, it fails for bounded MSD like those of the monomers relative to the ring CM, where after a suitably delayed time any of our dynamic processes would be predicted to independently contribute an MSD of $\lim_{t\rightarrow\infty}g_{2}(t)=2\langle R_{g}^{2}\rangle$.
To arrive at an alternative formulation, we note that for our definition of the relaxation time, $\tau_{X}=\langle R_{g}^{2}\rangle/D_{X}$, the null model suggests
$\frac{1}{\tau_{AB}}$ $=$ $\frac{p_{A}}{\tau_{A}}+\frac{p_{B}}{\tau_{B}}\ .$ (19)
Now consider
$\frac{1}{t_{AB}(g)}$ $=$ $\frac{p_{A}}{t_{A}(g)}+\frac{p_{B}}{t_{B}(g)}\ ,$ (20)
where $t_{\text{X}}(g)=g_{\text{X}}^{-1}(t)$ denotes the inverse function of $g_{\text{X}}(t)$, i.e. the time $t$ required with dynamics $X$ to reach an MSD of $g$. $t_{AB}(g)$ defined by Eq. (20) and its inverse function $g_{AB}(t)$ are well-behaved in a number of limiting cases. For simple diffusion, $g_{X}(t)\sim D_{X}t$ so that $t_{X}(g)\sim g/D_{X}$, Eq. (20) reduces to Eq. (11). By construction, $\lim_{p_{B}\rightarrow 0}t_{AB}(g)=t_{A}(g)/p_{A}$: as accepting only a fraction $p_{A}$ of the $A$ moves has the effect to reduce the effectively elapsed time by a corresponding factor, the time $t_{AB}$ required to reach a given MSD of $g$ increases by a factor $1/p_{A}$ relative to the time $t_{A}$ required with the pure $A$ dynamics. In particular, we thus find for the inverse function $\lim_{p_{B}\rightarrow 0}g_{AB}(t)=g_{A}(p_{A}t)$ as required. Plausibly, if both processes reach $g$ equally fast, $t_{A}(g)=t_{B}(g)$, then Eq. (20) suggests that the time required by the mixed process
$t_{AB}(g)$ $=$ $\frac{t_{A}(g)\,t_{B}(g)}{p_{A}t_{B}(g)+p_{B}t_{A}(g)}$ (21)
reduces to $t_{AB}(g)=t_{A}(g)=t_{B}(g)$ for arbitrary $p_{A}+p_{B}=1$. In contrast, if one of the processes reaches $g$ much faster than the other, $t_{A}(g)\ll t_{B}(g)$, then over a wide range of $p_{A}$-values $t_{AB}(g)\approx\frac{t_{A}(g)}{p_{A}}$ is essentially given by the time required by the faster process divided by the probability to activate the corresponding moves in the combined dynamics. Note that this relation holds in particular, when the second process by itself never reaches $g$ (i.e. when $t_{B}(g)=\infty$) and that according to Eq. (20) the mixed process cannot reach MSD, which none of the pure processes can reach individually ($t_{AB}(g)=\infty$, if $t_{A}(g)=t_{B}(g)=\infty$).
@@
Appendix B Comparison of the null model to simulation data for double-folded rings
Figure 9: MSDs for ideal double-folded rings, resulting from the combination of slithering dynamics and quenched tree dynamics, as a function of the mixing fraction $p$ introduced in Eq. 10. Here, $p_{\text{s}}$ denotes the fraction of slithering move in one MC sweep. The dashed gray lines in panels (a), (b), and (c) validate the null model using Eq. 20. Panel (d): the CM diffusion constant for the mixed dynamics across various chain lengths nomalized by the diffusion constant pure QTree dynamics ($P_{slithering}=0$). Dashed lines represent the null model predictions for the corresponding chain lengths as described by Eq. 11.
Figure 10: MSDs resulting from the superposition of Quenhced Tree (QTree) Dynamics and Ring-in-an-array-of-obstacles-like (RiAO) dynamics, based on the mixing fraction $p_{m}$ introduced in Eq. 10. $p_{\text{m}}$ represents the fraction of monomer move in one MC sweep and allows to interpolate between pure Ring-in-an-array-of-obstacles-like (RiAO) dynamics ($p_{m}=0$) and pure Quenched Tree Dynamics ($p_{m}=1$). Colored data points represent simulation data, while the thin gray lines indicate the predictions of the null model, Eq. 20, based on the simulation data for the pure case with $p_{m}=0$ and $p_{m}=1$.
Conveniently Eq. (20) is almost trivial to implement from a practical point of view, since it is straightforward (i) to generate numerically evaluable functions $t(g)$ and $g(t)$ from data sets associating paired values $(t,g)$ using standard interpolation routines and (ii) to invert the position of the dependent and the independent variable in plotted graphs. We have tested Eq. (20) for a case where we expect the null model to work: the dynamics emerging from the combination of repton and tree node moves, which are completely decoupled within our model. Panels (a)-(c) in Fig. 9 show $g_{1}(t)$, $g_{2}(t)$, and $g_{3}(t)$ for different proportions of repton and node moves, with $0\leq p_{m}\leq 1$. Only accepted monomer moves are Repton moves, with $q_{rep}=1$, and $q_{hp}=0$ (for more details, see Sec. III.4). For all considered values of $p_{m}$ our simulation data (shown as colored symbols) are in excellent agreement with the lines representing corresponding estimates based on Eq. (20) and our data for pure slithering and pure Brownian Tree Dynamics. Admittedly a bit trivially in the present case where $D_{\mathrm{Slithering}}=0$, the measured long time diffusion constants shown in Panels (d) also verify Eq. (11).
Figure 10 illustrates the results of the same analysis in the physically interesting case of an emergent dynamics to which monomer and node moves contribute in different proportions, $0\leq p_{m}\leq 1$ Eq. 10, such as to interpolate between pure Ring-in-an-array-of-obstacles-like (RiAO) dynamics ($p_{m}=1$) and pure Brownian Tree Dynamics ($p_{m}=0$). For non-interacting ideal rings, the predictions of the null model are again in excellent agreement with the simulation data (top row panels). For isolated self-avoiding rings (middle row panels), there are small, but noticeable deviation. Importantly, the accelerated dynamics of double-folded rings in a melt (bottom row panels) is clearly not described by the null model.
@@ Appendix B Comparison of the null model to simulation data for double-folded rings
Appendix C The emergence of the ring-in-an-array-of-obstacles-like dynamics from the combination of hairpin and repton moves
Figure 11: The MSDs when modifying the relative weight of repton and hairpin move in the ring-in-an-array-of-obstacles-like (RiAO) dynamics. The Inset panel shows the local slopes of the data, and the horizontal dashed lines depict the expected behavior of RiAO dynamics. In Panel (D), the symbols represent CM diffusion constant.
To better understand possible signatures of non-trivial emergent dynamics arising from the interplay of two different local moves, we have modified the relative weight of Repton and Hairpin moves in the ring-in-an-array-of-obstacles-like dynamics (Fig. 11). Since neither Repton nor Hairpin moves are ergodic by themselves, the null model clearly does not apply: $D_{\text{Rep}}=D_{\text{HP}}=0$, Eq. (19) would predict $D_{\text{RiAO}}=0$ And yet the rings move and equilibrate as long as both types of moves are present.
In particular, we observe the characteristic anomalous monomer and CM diffusion predicted by the ring-in-an-array-of-obstacles model equally well, when 99% of the Repton moves are suppressed as when 99% of the Hairpin moves are rejected (Panels (a)-(c) in Fig. 11). There is, however, a large quantitative difference in that suppressing 99% of the Hairpin moves has only a small effect on the relaxation time, while suppressing 99% of the Repton moves slows down the overall dynamics by a corresponding factor.
To analyse this situation, we denote as $f_{A}$ and $f_{B}$ the fractions of accepted $A$ (Repton) and $B$ (Hairpin) type MC moves per sweep or unit of time; the remaining moves are rejected and do not contribute to the dynamics. Now suppose that $a$ Repton moves and $b$ Hairpin moves are required for an effective step of the ring-in-an-array-of-obstacles-like dynamics. The resulting fraction of RiAO moves,
$f_{aA+bB}=\text{min}\left(\frac{f_{A}}{a},\frac{f_{B}}{b}\right)\ ,$ (22)
generalizes to
$f_{aA+bB}=\text{min}\left(\frac{q_{A}f_{A}}{a},\frac{q_{B}f_{B}}{b}\right)\ ,$ (23)
if one changes the relative weight of Repton and Hairpin moves by accepting them only with reduced probabilities $0\leq q_{A},q_{B}\leq 1$.
Importantly, $f_{aA+bB}$ is proportional to the asymptotic CM diffusion constant for the emergent coupled dynamics: $D_{aA+bB}\sim f_{aA+bB}$. For our choice of $q_{A}=q$ and $q_{B}=1-q$, Eq. (23) takes the simple form of two straight lines
$f_{aA+bB}=\left\{\begin{array}[]{cc}q\,\frac{f_{A}}{a}&\mathrm{if}\ q<q^{\ast}\\ (1-q)\frac{f_{B}}{b}&\mathrm{if}\ q>q^{\ast}\end{array}\right.$ (24)
originating at $f_{aA+bB}\equiv 0$ for $q=0$ and $q=1$ and intersecting at $q^{\ast}=af_{B}/(bf_{A}+af_{B})$ where $f_{aA+bB}(q)$ adopts its maximum value of $f_{aA+bB}(q^{\ast})=f_{A}f_{B}/(bf_{A}+af_{B})$.
The prediction is in good qualitative agreement with our simulation results shown in Fig. 11(d). The interpretation is straightforward. The existence of a maximum at finite $10<q=q^{\ast}<$ 1 is indicative of a dynamics arising from a combination of moves. At the maximum $\frac{q_{B}f_{B}}{b}=\frac{q_{A}f_{A}}{a}$, so that Repton and Hairpin moves are attempted in “stochiometric” proportions. $q>q^{\ast}$ corresponds to $\frac{q_{B}f_{B}}{b}<\frac{q_{A}f_{A}}{a}$, where the limiting factor is the fraction of successful Hairpin moves. In the opposite case of $q<q^{\ast}$ and $\frac{q_{B}f_{B}}{b}>\frac{q_{A}f_{A}}{a}$ the combined dynamics is limited by the fraction of accepted Repton moves. The latter case corresponds to our simulations, where the vast majority of attempted monomer moves correspond to hairpin moves. In particular, the above analysis explains, why the frequent rejection of hairpin moves in our melts has such a small effect on the amplitude of the RiAO dynamics (Figure 4).
The effect of combining Repton and Hairpin moves thus seems at the same time spectacular and trivial. It is spectacular in that the emergent ring-in-an-array-of-obstacles-like dynamics is qualitatively completely different from what one observes in the two non-ergodic limiting cases of pure Hairpin or Repton moves: The rings become mobile and are able to anneal the connectivity of the trees characterizing the double-folding. But the way how the two moves combine appears to be as trivial as for the combination of hopping moves in the $x$- and $y$-directions on a square lattice in the case of a single particle diffusing along a tilted narrow strip. In particular, the particle remains localized, if it attempts to hop exclusively along one or the other of the axes of the lattice. And as in the case of our Repton and Hairpin moves, there is an optimal mixture of $x$- and $y$-moves, which in this example trivially depends on the orientation of the strip relative to the lattice.

### v2

@@ VI.  SUMMARY AND CONCLUSION
Ring polymers pose a number of challenging fundamental problems in polymer physics 25, 31, 32. In this article, we have studied the dynamics of a (lattice) model of tightly double-folded chains 1, 2, 3, 6, 38. In our Monte Carlo simulations, ring monomer moves are responsible for the spontaneous creation and deletion of side branches and the diffusive local mass transport along the tree, while the exclusive application of tree node moves gives rise to a Brownian dynamics of double-folded rings with quenched secondary structure. In particular, the present generalization of our earlier model 38 allows us to explore arbitrary combinations of ring and tree dynamics where sets of compatible ring and tree degrees of freedom are simultaneously updated and defined at all moments of time.
@@
As a starting point, we have explored the dynamics emerging from different combinations of local moves for three paradigmatic types of tree-like double-folded ring systems: ideal rings, isolated self-avoiding rings, and rings in the melt state. The local transport of reptons gives rise to a slithering of rings around a fixed tree, while the ring-in-an-array-of-obstacles (RiAO) dynamics emerges in the presence of additional monomer moves, which allow to absorb and grow side branches. The QTree dynamics of quenched trees undergoing Brownian motion turns out to be Rouse-like for ideal and isolated self-avoiding rings. In the melt state, the monomer diffusion is again in good agreement with the expectation $g_{1}\sim t^{0.4}$ from the Rouse model for crumpled rings. However, there are also clear indications for a slowdown of the center-of-mass dynamics on intermediate scales, which translates into a reduced center-of-mass diffusion $D_{\mathrm{QTree}}\sim N^{-1.6\pm 0.1}$.
To apply the model to a particular ring polymer system (see e.g. 9, 13, 14, 15, 16, 18, 19, 20), one needs to specify the relative importance of the different types of local dynamics represented by the ring monomer and tree node moves. For this purpose we have introduced a control parameter ($0\leq p_{m}\leq 1$) allowing us to explore the full range of dynamical scenarios, where the limiting cases are RiAO ($p_{m}=1$) and QTree ($p_{m}=0$) dynamics. The appropriate value of $p_{m}$ for a physical system reflects the microscopic mobility of individual monomers relative to the collective mobility of tree-like segments. These mobilities depend on several physical factors, including local friction, chain stiffness, and topological constraints. For example, in RNA molecules, the secondary structure typically corresponds to a quenched tree-like system. In contrast, dense polymer melts may be better described by RiAO dynamics. For small, supercoiled DNA molecules, an intermediate value of $p_{m}$ is more likely, as both slithering dynamics 59 and the creation/annihilation of branches 60, 61 contribute to the overall behavior. By comparing MSDs measured in physical or biological systems, one should be able to infer the value of $p_{m}$ that best represents the dynamics at play.
Our objective here was to study generic aspects of the coupling of different local modes of dynamics. In particular, we have exploited the opportunity inherent in a mesoscale description to freely vary the relative kinetic rates instead of being limited to ratios emerging in particular experimentally or numerically accessible target or model systems. For non-interacting ideal rings we found that the dynamics emerging from the simultaneous presence of monomer and node moves are well described by an additive null model. In contrast, a non-trivial acceleration emerges for interacting rings, which we have analyzed in detail in Sec. 5. For self-avoiding trees, the corresponding contribution to the ring center-of-mass diffusion constant scales with the ring size like $\delta D\sim N^{-1}$. This corresponds to a sub-dominant acceleration of the Rouse-like QTree dynamics. In the melt case, the coupling contribution dominates. While the scaling is less clear, our data are compatible with the interpretation that constraint release due to the annealing of the tree structure asymptotically induces a Rouse-like dynamics.
Finally, we note that numerous experimental investigations using fluorescence microscopy have explored the 3D dynamics of chromatin in the cell nuclei of mammals 62, Drosophila 63, 64, and yeast 63, 65 as well as in the bacterial nucleoid 66, 67, 68, 69. Reported results show a variety of subdiffusive behaviors, with $g_{1}$ exponents typically ranging from $\sim 0.3$ to $\sim 0.7$, depending on cell activity, environment, and the specific chromatin locus under study. In this context, our observation of an exponent $\simeq 0.4$ in the dynamics of monomers in dense ring melts is strikingly reminiscent of the typical exponent of $0.4$ observed in bacterial chromosomes 67, 68, 69. While several scenarios have been proposed to explain this value (see, e.g. 70 and references therein), it is notable that a tree-like, double-folded ring corresponds to the typical structure associated with supercoiled DNA, a hallmark of bacterial chromosomes 9, 13. Nonetheless, this correspondence should be interpreted with care. Our study concerns Brownian-like, near-equilibrium situations where the only source of energy is the thermal activity of the (implicit) surrounding medium. In contrast, chromatin dynamics in vivo is shaped by many active, far-from-equilibrium processes such as ATP-dependent remodeling, transcriptional activity, DNA replication, and loop extrusion 71, 13, 72. Additionally, chromatin exhibits structural and topological complexity that is not captured by our models, and the relevance of our results may depend on the time and length scales considered. In this regard, observations in eukaryotes 73, 74 of exponents in the range of 0.50.60.5-0.6 beyond the entanglement or branching scale call for different explanations than those advocated in the present work.
VII.  SIMULATION SOFTWARE
@@
The simulation software is an open-source package, distributed under the GNU General Public License v3.03.0 and written in C++. It is available at 75. All simulations were performed on CBPSmn computer cluster of ENS-Lyon.
VIII.  ACKNOWLEDGMENT
@@
We gratefully acknowledge the support of the Centre Blaise Pascal’s IT test platform at ENS de Lyon (Lyon, France) for facilities. The platform operates the SIDUS solution 76 developed by Emmanuel Quemener. We thank Angelo Rosa and Cedric Vaillant for valuable discussions, and Ali Farnudi for his technical support. IJ and RE acknowledge funding from CNRS 80 Prime (MIMIC project). RE has greatly benefited from discussions with Michael Rubinstein, Dimitris Vlassopoulos and Charles M. Schroeder at and after the 2023 Cecam workshop on ring dynamics in Prato.
IX.  SUPPORTING INFORMATION
Supporting Information includes:
  • Sec. I: Equilibration
  • Sec. II: Effect of maximal functionality of tree nodes on the dynamics
  • Sec. III: Stored-Length (Repton) Distribution
@@ IX.  SUPPORTING INFORMATION
  • Sec. IV: Approximate solution of the null model for the dynamics resulting from independent additive contributions
@@
  • Sec. V: Comparison of the null model to simulation data for double-folded rings
@@
  • Sec. VI: The emergence of the ring-in-an-array-of-obstacles-like (RiAO) dynamics from the combination of hairpin and repton moves
  • Sec. VII: Possible contributions to the ring CM diffusion constant from two coupled dynamics
  • Sec. VIII: Brownian dynamics of randomly quenched trees (QTree dynamics)
  • Sec. IX: Ring-in-an-array-of-obstacles-like (RiAO)dynamics
  • Table I: Summary of system parameters for polymer chains.
  • Table II: Diffusion constants $D$ for different systems, sizes, and mixing ratios.
  • Figure S1: Equilibration analysis.
  • Figures S2-S3: Effect of maximum node functionality on structural and dynamical properties.
  • Figures S4: Distribution of tree node functionalities for systems with different imposed maximal functionalities ($f_{max}=3,6,9,12$).
  • Figure S5: Stored-length (repton) distribution and Poisson model.
  • Figure S6: MSDs for ideal double-folded rings, resulting from the combination of slithering dynamics and quenched tree dynamics, as a function of the mixing fraction $p_{m}$.
  • Figure S7: MSDs resulting from the superposition of Quenhced Tree (QTree) dynamics and the Ring-in-an-array-of-obstacles-like (RiAO) dynamics, based on the mixing fraction $p_{m}$.
  • Figure S8: The MSDs and the difussion constants for varying the relative weights of repton and hairpin move in the ring-in-an-array-of-obstacles-like (RiAO) dynamics.
  • Figures S9–S11: Brownian quenched tree dynamics (Ideal, Self-avoiding, Melt systems).
  • Figures S12–S14: Ring-in-an-array-of-obstacles-like (RiAO) dynamics (Ideal, Self-avoiding, Melt systems).
  • Figure S15: Slithering dynamics of ideal rings.
  • Figure S16: Scaling of the CM diffusion constant $D$ for varying mixing ratios $p_{m}$.
  • Figure S17: Relaxation times and CM diffusion constants versus $p_{m}$ compared to null model.